# A Simple Knowledge Distillation Framework for Open-world Object Detection

### v1

@@ Conclusions
In this paper, we attempt to specialize the VLM model for OWOD task by distilling its open-world knowledge into a language agnostic detect. we observe that the combination of the simple knowledge distillation and automatic pseudo-labeling mechanism could convert open-world knowledge. To mitigate the effect of distilled knowledge on the detection performance of known objects, we propose the down-weight training loss function for the detector’s mixed learning process of known ground truth and distilled unknown knowledge. We propose a cascade decoupled detection transformer structure to alleviate the influence caused by unknown objects on detecting known objects. Our extensive experiments on two popular benchmarks demonstrate the effectiveness of our framework.
Acknowledgment
@@
This work is supported by National Nature Science Foundation of China (grant No.61871106), and the Open Project Program Foundation of the Key Laboratory of Opto-Electronics Information Processing, Chinese Academy of Sciences (OEIP-O-202002).

### v2

@@ 8 Conclusions
In this paper, we start by observing that the simple knowledge distillation could convert the open-world knowledge in the large pre-trained vision-language grounding model for the specialized OWOD task and propose a simple framework with surprisingly good performance. Meanwhile, we propose the down-weight training loss function for the detector’s mixed learning process of known ground truth, distilled unknown knowledge, and the pseudo unknown knowledge in OWOD algorithm to mitigate the effect of distilled knowledge on the detection performance of known objects. Besides, the cascade decoupled detection transformer structure is proposed to alleviate the influence caused by unknown objects on detecting known objects. Last but not least, we propose two novel benchmarks to comprehensively evaluate the ability of the open-world detectors to detect unknown open-world objects, named StandardSet and IntensiveSet respectively, based on the complexity of their testing scenarios. Our extensive experiments on existing and proposed benchmarks demonstrate the effectiveness of our framework. Our model exceeds the distilled large pre-trained vision-language grounding model for OWOD and state-of-the-art methods for OWOD and IOD.
Acknowledgments
@@
This work is supported by National Natural Science Foundation of China (grant No.61871106 and No.61370152), Key R&D projects of Liaoning Province, China (grant No.2020JH2/10100029), and the Open Project Program Foundation of the Key Laboratory of Opto-Electronics Information Processing, Chinese Academy of Sciences (OEIP-O-202002).